‏Suppose a laser starting an original field strength Eo bounces back and forth between two mirrors with reflectivity R1 and R2 . The field makes a round trip in a time T = 2L / v , where v is velocity . The electric field will decay with time and the decaying field will be?


1
Expert's answer
2022-04-24T17:18:57-0400

"E=\frac{E_0}{1-e^{-kT}}\cdot \frac{R_2}{R_1}."